OpenStudy (briensmarandache):

solve for x: Sqrt(1+cosX) - Sqrt(2)cos(x/2) = Sqrt(2) the radical 2 is not including cos x/2

OpenStudy (briensmarandache):

@kropot72

OpenStudy (briensmarandache):

@perl would you be able to help me with this?

OpenStudy (perl):

add sqrt(2)cos(x/2) to both sides

OpenStudy (briensmarandache):

got that, but not sure where to go from there after i square both sides

OpenStudy (briensmarandache):

@perl not seeing how to continue

OpenStudy (perl):

Sqrt(1+cosX) - Sqrt(2)cos(x/2) = Sqrt(2) sqrt(1 + cos x ) = sqrt(2)cos(x/2) + sqrt(2) square both sides 1 + cos x = 2cos^2(x/2) + 2 sqrt(2)cos(x/2) * sqrt(2) + sqrt(2)^2

OpenStudy (briensmarandache):

hold on this format is a little confusing

OpenStudy (perl):

actually theres a simpler way to do this , i think

OpenStudy (perl):

notice that sqrt( 1 + cos x) , is very close to sqrt ( (1 + cos x ) / 2)) , which is equal to cos(x/2)

OpenStudy (perl):

Sqrt(1+cosX) - Sqrt(2)cos(x/2) = Sqrt(2) divide both sides of equation by sqrt(2)

OpenStudy (perl):

unfortunately my draw box does not work , i get invisible ink

OpenStudy (briensmarandache):

not seeing how the previous is similar to the latter, is this a rule or something?

OpenStudy (perl):

there is a trig identity called half angle identity http://www.purplemath.com/modules/idents.htm cos(x/2) = sqrt( (1 + cos x ) / 2)

OpenStudy (perl):

so our equation has a pretty close expression

OpenStudy (perl):

watch what happens when you divide both sides by sqrt(2)